The graph to the right shows a region of feasible solutions. Use this region to find maximum and minimum values of the given objective functions, and the locations of these values on the graph. 10- (2, 8) (a) z -6x +9,y (b)2-x + 3y (a) What is the maximum of z 6x+ 9y? Select the correct answer below and, if necessary, fill in the answer boxes to complete your choice. (7, 5) (0 A. The maximum value of the objective function z- 6x + 9y is 87 and is located at the point (7,5) Type an exact answer.) 10 O B. The maximum does not exist. What is the minimum of z 6x+9y? Select the correct answer below and, if necessary, fill in the answer boxes to complete your choice. O A. The minimum value of the objective function z- 6x + 9y is located at the point (Type an exact answer.) and is O B. The minimum does not exist.
Answer #1

2nd question:

A.the minimum value of the objective function Z=6x+9y is 27 and is located at the point (3,1).

working :

at (3,1) the value will be = 6*(3) + 9*(1) =>27.

at (0,6) the value will be = 6*(0) + 9*(6)=>54.

at (2,8) the value will be = 6*(2) + 9*(8)=>84.

the lowest value is 27 at the point (3,1).